OpenStudy (anonymous):

Simplify: (3x+2y)(3x-2y)=

jimthompson5910 (jim_thompson5910):

this is a difference of squares use the difference of squares formula \[\Large (A+B)(A-B)=A^2-B^2\]

OpenStudy (anonymous):

where A=3x and B=2y

OpenStudy (anonymous):

so is it 9x2-4y2

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

\[=(3x)^2-(2y)^2=9x^2-4y^2\]

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

dang u guys are so fast lol... wow i really need this sites help... i been taking algebra 1 four times cuz im terrible at math... im genius in other subjects but math is hard for me.

OpenStudy (anonymous):

and im in my second yr as a senior... and i need four math credits to graduate

OpenStudy (anonymous):

Just in case you don't know about it - you might want to kick back , relax and enjoy some algebra videos on www.khanacademy.org focusing on topics which are giving you grief. There are also online practice questions available.

OpenStudy (anonymous):

thanks...

OpenStudy (anonymous):

thanks...

OpenStudy (anonymous):

so can someone help this girl ? -.-